# Raising the Bar of AI-generated Image Detection with CLIP

### v1

@@ 4.1 Influence of the reference set size
We begin with investigating the role of the most important parameter of the proposed detector, the number of real / synthetic samples $N$ in the reference set, that in this experiment includes COCO real images and Latent Diffusion generated ones. Fig. 3 shows results for $N$ varying from 10 to 100k in log scale. The results are presented in terms of AUC, AP, and Accuracy. When $N$ is less than 10k, we perform several runs with different sets of images and report the average metrics, e.g. for $N$ equals to 1k, we show the average on 10 runs.
Let us focus first on the top row, showing results on images that have been neither re-compressed nor resized. In this simpler scenario, AUC and AP are extremely high, always above 85% for all known families of synthetic images (GAN, Diffusion) and somewhat worse for images generated by commercial tools. Especially remarkable is the fact that very good results are obtained already with as few as 10+10 reference images, showing this lightweight solution to be fully viable. The performance improves with $N$, growing by up to 5-10 percent points at the $N=10k$ plateau. We observe, in passing, that AUC and AP curves are very similar, as always happens when they approach 100%, hence we drop the latter for brevity in further experiments.
@@
In the rightmost figure, we see that unlike AUC and AP, accuracy does not always improve with increasing $N$. This is because AUC and AP are integral threshold-independent metrics. They tell us how well a method could do with a perfectly calibrated threshold. On the contrary, accuracy depends critically on the selected threshold which, lacking any prior knowledge, is set to 0.5. For Latent and GAN, 0.5 is a good threshold, and accuracy improves as the classifier does. Instead, for Diffusion and Commercial, the optimal threshold grows farther from 0.5 as $N$ increases and the accuracy decreases significantly. So there is a trade-off between optimality and robustness in our truly OOD scenario, where no calibration data exist (it is very unlikely that we know exactly which is the generator of the data under test). On the other hand, if we had data from the target class, even just 10 images, we could use them to design and ad hoc classifier.
The bottom row of the figure shows results for images that have been compressed and/or resized. As expected, performance degrades somewhat in this scenario, but continues to be very good, with AUC between 75% and 90% and accuracy between 65% and 80% at $N$=10k. These are excellent results, much beyond the current state of the art, as we will show later. In this case, using only 10+10 images seems not to be advisable but the performance is near-optimal already at $N$=100.
Datasets
augm.
GAN
Diffusion
Commerc.
AVG
family
family
tools
COCO $+$ Latent
93.9
93.3
81.7
89.6
COCO $+$ Latent
84.8
91.0
87.6
87.8
COCO $+$ ProGAN
95.0
90.0
65.6
83.5
COCO $+$ ProGAN
92.3
90.1
82.3
88.2
LSUN $+$ Latent
87.0
82.7
65.6
78.5
LSUN $+$ Latent
86.8
79.9
80.3
82.4
LSUN $+$ ProGAN
93.8
79.2
44.7
72.6
LSUN $+$ ProGAN
95.7
82.7
67.5
82.0
Table 2: Performance of the CLIP-based detector for various real/fake combinations, with and without augmentation.
@@ 4.2 Influence of the reference set content
We discovered that not only the size but also the quality of the reference set images impacts heavily on performance. This probably does not need proof if we think about a very small reference set, like 10+10 images, where low quality or low diversity would spell catastrophe. However, something similar happens also when a larger reference set is used. Tab. 2 shows the results in terms of AUC obtained using various combination of real and synthetic images in the reference set, that is, (Real, Synth) $\in\{{\rm COCO,LSUN}\}\times\{{\rm Latent,ProGAN}\}$, with and without augmentation. In all cases, the reference set has size 10000+10000. Results are clearly affected by the specific combination. Indeed, there is a marked impairment when the LSUN dataset is used instead of COCO to draw the real samples and also, to a lesser extent, when ProGAN is used instead of Latent for the synthetic images. We conjecture that both LSUN and ProGAN are not diverse enough to describe adequately the decision domain. Moreover, LSUN has several polarizations not included in COCO, the images of LSUN represent only few well-defined categories, have all the same size, and are compressed most of them with the same quality factor. In Tab. 2, we see that data augmentation only marginally reduces the effects of the used dataset.
@@ 4.3 Influence of pre-training
We experimented also with various versions of CLIP ViT L/14 [35], trained on different datasets, observing significant changes in performance. Fig. 4 shows, the AUC, AP and Accuracy averaged on all families of models as a function of the pre-training dataset. Each curve refers to a different value of $N$ and we consider only the case of post processed images for brevity. For both metrics and all values of $N$ there is a steady increase in performance as larger and larger datasets are used for pre-training, from LAION-400M (0.4B images) [64], to LAION (2B) [65], to CommonPool (12.8B) [32]. Overall, from smallest to largest dataset there is a ten point improvement! This comes to confirm, if ever necessary, the importance of pre-training a large VLM on the largest possible collection of varied images to acquire greater descriptive power for all kinds of applications.
4.4 Beyond low-level forensic traces
As already highlighted in the Introduction, forensic classifiers typically rely on subtle low-level artifacts. These can be seen more clearly in the frequency domain, where they appear as strong peaks. In this subsection we report on some initial analysis aimed at understanding whether the CLIP-based detector also relies on this type of traces or it exploits different information. To this end, we consider the high resolution 1024$\times$1024-pixel images generated by Stable Diffusion XL and resize them to 256$\times$256. Fig. 5 shows, on the right, the high and low-resolution images (top) together with their spectra (bottom). We observe that after resizing, the spectral artifacts have completely disappeared. In fact, these traces are caused by the aliasing effect of the upsampling filters, which can be removed by proper decimation. We run our detector on 1000 original and resized synthetic images, obtaining in both cases the same True Positive Rate (TPR) of $43\%$ (note that the True Negative Rate (TNR) for these images is $97\%$).
We also carry out a complementary experiment where we take real images and pass them through the generative architecture of Stable Diffusion which acts as an autoencoder. Therefore, the output (Fig. 5 right) is visually identical to the input but its Fourier spectrum now shows strong peaks. Even so, the CLIP-based detector is not affected by the presence of such traces and the TNR, $97\%$, is almost the same as with unprocessed images. We find this last experiment to be particularly relevant. In fact, it is very important to take into account malicious operations that can be carried out with the aim of reducing the credibility of current forensic detectors. Of course, several forms of attack can be applied, but this one seems to be especially simple and disruptive, such to easily fool detectors that rely exclusively on low-level traces. Based on this experiment, it seems that the proposed CLIP-based detector looks for different types of traces. Therefore, besides its stand-alone performance, it could be used to complement and enhance the performance of other detectors. A larger set of experiments can be found in the supplementary.

### v2

@@ 4.1 Influence of the reference set size
We start by investigating the role of the most important parameter of the proposed detector, the number of real / synthetic samples $N$ in the reference set, which in this experiment includes COCO real images and Latent Diffusion synthetic images. Fig. 3 shows results for $N$ ranging from 10 to 100k in log scale. Results are in terms of AUC, AP and Accuracy. When $N$ is less than 10k, we perform several runs with different sets of images and report the average metrics, e.g., for $N$ equals 1k, we average on 10 runs.
Let’s focus first on the top row, which shows results on images that were neither recompressed nor resized. In this simpler scenario, both AUC and AP values are extremely high, consistently above 85% for all known families of synthetic images (GAN, Diffusion) and significantly worse only for images generated by commercial tools. It’s particularly remarkable that very good results are achieved with as few as 10+10 reference images, proving that this lightweight solution is fully viable. The performance improves with $N$, growing up to 5-10% at the 10k plateau. We note, in passing, that AUC and AP curves are very similar, so we drop the latter for brevity in further experiments.
@@
In the rightmost figure, we see that, unlike AUC and AP, accuracy does not always improve as $N$ increases. This is because AUC and AP are threshold-independent integral metrics. They tell us how well a method with a perfectly calibrated threshold might work. Instead, the precision depends critically on the selected threshold which, in the absence of any prior knowledge, is set at 0.5. For Latent, 0.5 is a good threshold and accuracy improves as $N$ grows. Instead, in general, the optimal threshold moves further away from 0.5 as $N$ increases and the precision decreases significantly. So, there is a trade-off between optimality and robustness in our truly OOD scenario, where no calibration data exists. On the other hand, if we had data from the target class, even just 10 images, we could use them to design an ad hoc classifier (see supplementary material).
The bottom row of the figure shows results for images that have been compressed and/or resized. As expected, performance degrades slightly in this scenario, but continues to be very good, with AUC between 75% and 90% and accuracy between 65% and 80% at $N$=10k. These are excellent results, well beyond the current state of the art, as we will show later. In this case, using only 10+10 images does not seem advisable but the performance is almost optimal already at $N$=100.
Datasets augm. GAN Diffusion Commerc. AVG
family family tools
COCO $+$ Latent 92.4 92.6 80.5 88.5
COCO $+$ Latent 89.3 91.8 87.0 89.4
COCO $+$ ProGAN 93.6 90.0 65.6 83.1
COCO $+$ ProGAN 91.3 90.1 82.3 87.9
LSUN $+$ Latent 88.7 82.7 65.6 79.0
LSUN $+$ Latent 87.9 79.9 80.3 82.7
LSUN $+$ ProGAN 94.1 79.2 44.7 72.7
LSUN $+$ ProGAN 95.0 82.7 67.5 81.7
Table 2: AUC performance of the CLIP-based detector for various real/fake data, with and w/o augmentation (resizing/compression).
@@ 4.2 Influence of the reference set content
We found that not only the quantity but also the quality of images in the reference set significantly impacts performance. This is intuitive with a very small reference set, like 10+10 images, where low quality or limited diversity could have catastrophic consequences. However, similar effects are observed even with a larger reference set. Tab. 2 shows the results in terms of AUC obtained using various combination of real and synthetic images in the reference set, that is, (Real, Synth) $\in\{{\rm COCO,LSUN}\}\times\{{\rm Latent,ProGAN}\}$, with and without augmentation. In all cases, the reference set has size 10,000+10,000. Results are clearly influenced by the specific combination. Indeed, considerable degradation occurs when the LSUN dataset is used instead of COCO to draw the real samples and, to a lesser extent, when using ProGAN instead of Latent for the synthetic images. We conjecture that both LSUN and ProGAN lack the diversity necessary to adequately describe the decision domain. Furthermore, unlike COCO, the LSUN dataset exhibits several biases: the images represent only a few well-defined categories, have all the same size, and most of them are compressed with the same quality factor. In Tab. 2, we see that data augmentation only marginally reduces the effects of the dataset used.
@@ 4.3 Influence of pre-training
We experimented also with various versions of CLIP ViT L/14 [36], trained on different datasets, observing significant changes in performance. Fig. 4 shows, the AUC, AP and Accuracy averaged on all families of models as a function of the pre-training dataset. Each curve refers to a different value of $N$ and we consider only the case of post processed images for brevity. For both metrics and all values of $N$ there is a steady increase in performance as increasingly larger datasets are used for pre-training, from LAION-400M (0.4B images) [65], to LAION (2B) [66], to CommonPool (12.8B) [33]. Overall, there is a ten-point improvement from the smallest to the largest dataset. This confirms the importance of pre-training a large VLM on the largest possible collection of different images.
Method Real/Synth. Training Size (k) Aug. Test Strategy
[73] Wang et al. LSUN / ProGAN 360 / 360 global pooling
[9] PatchFor. CelebA,FF / various 84 / 272 resizing
[34] Grag. et al. LSUN / ProGAN 360 / 360 global pooling
[49] Mand. et al. various / various 232 / 386 patch aggregation
[48] Liu et al. LSUN / ProGAN 360 / 360 global pooling
[14] Corvi et al. COCO,LSUN / Latent 180 / 180 global pooling
[70] LGrad LSUN / ProGAN 72 / 72 resizing
[57] Ojha et al. LSUN / ProGAN 360 / 360 cropping
[74] DIRE-1 LSUN-Bed / ADM 40 / 40 resizing
[74] DIRE-2 LSUN-Bed / St.GAN 40 / 40 resizing
[71] NPR LSUN / ProGAN 72 / 72 resizing
Ours 1k COCO / Latent 1 / 1 resizing
Ours 1k+ COCO / Latent 1 / 1 resizing
Ours 10k COCO / Latent 10 / 10 resizing
Ours 10k+ COCO / Latent 10 / 10 resizing
Table 3: List of methods. For each method, we report the datasets of real and synthetic images used for training, their sizes, whether or not augmentation is used, and the testing strategy.
GAN family Diffusion family Commercial tools AVG
Pro Style Style Style Giga Score ADM GLIDE Latent Stable DeepFl. Ediff-I DiT SDXL DALL· DALL· Midj. Adobe
Method GAN GAN2 GAN3 GANT GAN SDE Diff. Diff. IF E2 E3 Firef.
Wang et al. 100. 96.5 98.5 98.9 66.6 32.9 64.3 48.5 59.2 41.5 78.0 64.9 58.6 54.3 64.8 10.9 40.2 84.8 64.6
PatchFor. 92.3 84.5 91.8 91.2 64.7 83.3 74.8 96.2 78.1 62.4 62.7 78.7 83.1 68.4 41.9 52.7 57.8 49.4 73.0
Grag. et al. 100. 99.8 97.5 98.8 82.8 92.1 74.7 62.8 91.9 52.5 69.9 69.6 65.3 58.0 58.3  2.4 43.1 63.5 71.3
Mand. et al. 96.2 93.8 100. 92.6 61.8 99.8 56.5 40.5 70.0 36.8 47.2 65.0 59.1 27.0 14.5 14.7 24.3 36.7 57.6
Liu et al. 100. 99.8 98.4 98.5 98.2 95.4 82.5 76.5 97.6 77.4 72.2 98.7 88.0 31.1 70.4  0.2 40.7 11.8 74.3
Corvi et al. 79.4 73.7 50.0 97.1 63.4 65.0 80.7 91.9 100. 100. 99.9 85.7 100. 100. 69.4 60.8 100. 98.0 84.2
LGrad 100. 91.2 83.8 81.8 82.2 80.6 76.9 66.1 81.1 61.5 68.8 74.1 56.2 57.2 58.6 37.9 56.3 40.6 69.7
Ojha et al. 100. 93.9 92.3 98.2 96.0 58.4 86.7 80.8 85.7 89.5 92.9 80.6 77.8 85.1 95.2 36.4 66.2 97.5 84.1
DIRE-1 50.6 56.9 47.8 99.9 74.1 44.3 75.7 71.4 68.7 39.4 98.9 99.1 99.6 47.1 44.7 47.6 51.0 57.4 65.2
DIRE-2 54.2 52.5 43.0 99.6 76.0 41.0 70.1 70.1 69.3 46.9 97.0 98.2 98.3 42.8 41.0 49.6 47.8 43.0 63.3
NPR 100. 85.6 77.0 96.4 88.7 91.1 86.3 79.3 90.2 64.5 91.6 80.1 78.4 76.7 39.5 48.7 77.0 32.1 76.8
Ours 1k 98.9 90.5 85.5 100. 81.3 89.1 81.1 99.9 94.1 87.6 96.5 98.5 94.1 87.8 89.0 70.0 73.0 74.4 88.4
Ours 1k+ 91.4 80.9 84.0 99.8 74.7 84.3 75.2 99.6 81.6 89.8 98.0 99.1 92.5 88.9 83.6 93.6 78.7 85.1 87.8
Ours 10k 99.8 91.8 86.8 100. 83.6 89.0 81.4 99.9 94.2 90.7 97.0 98.7 95.0 87.4 89.2 77.6 75.3 80.1 89.8
Ours 10k+ 93.4 87.1 87.6 99.9 78.5 89.2 79.9 99.7 84.7 91.3 97.9 99.4 94.0 90.1 86.3 92.9 81.7 87.2 90.0
Table 4: Comparison with SoTA methods in terms of AUC. For our approach we show four variants: 1k and 10k indicate the number of real and fake training images, + indicates augmentation (compression/resizing). Results on the dataset used for training are in light gray, while bold underlines the best performance for each dataset with a margin of 1%. The last column shows the average over all datasets.
GAN family Diffusion family Commercial tools AVG
Pro Style Style Style Giga Score ADM GLIDE Latent Stable DeepFl. Ediff-I DiT SDXL DALL· DALL· Midj. Adobe
Method GAN GAN2 GAN3 GANT GAN SDE Diff. Diff. IF E2 E3 Firef.
Wang et al. 100. 86.6 88.4 61.7 59.2 68.0 65.0 60.6 67.1 55.2 50.3 48.3 55.1 64.5 46.2 27.7 46.7 55.9 61.5
PatchFor. 57.9 51.8 57.0 50.6 53.8 69.0 66.2 83.3 58.8 48.3 61.3 65.0 68.1 63.3 64.3 63.3 59.0 65.1 61.4
Grag. et al. 100. 95.4 90.9 94.4 64.4 77.1 77.1 79.8 84.0 53.5 50.6 55.6 66.7 66.6 55.2 25.1 48.5 60.2 69.2
Mand. et al. 81.1 79.3 87.2 49.1 49.3 64.0 54.8 42.6 52.9 39.4 55.7 54.5 49.8 42.2 47.9 42.3 35.2 53.4 54.5
Liu et al. 64.3 55.1 50.1 57.3 45.4 62.6 51.1 58.6 50.7 58.6 50.9 64.2 53.9 56.0 44.4 61.8 52.6 53.1 55.0
Corvi et al. 77.5 74.7 69.4 82.1 66.6 70.4 79.0 93.5 99.3 69.9 60.7 72.1 89.2 61.8 65.9 32.4 51.9 58.1 70.8
LGrad 56.3 58.3 49.8 52.3 43.5 45.9 49.2 42.3 50.4 54.8 40.7 46.4 49.4 53.2 41.8 53.5 50.4 51.8 49.4
Ojha et al. 99.8 75.5 75.4 91.1 88.5 79.3 83.7 83.3 81.8 75.0 59.9 68.7 70.1 61.8 63.2 41.7 40.6 52.9 71.8
DIRE-1 48.4 42.5 39.1 53.5 54.3 44.1 48.0 44.7 46.1 47.0 66.2 62.8 53.2 47.1 44.6 47.6 51.0 57.4 49.9
DIRE-2 49.3 41.6 38.6 53.8 55.0 44.3 45.1 40.2 45.9 56.4 70.7 72.2 53.0 42.8 40.9 49.7 47.8 43.0 49.5
NPR 54.5 48.5 41.9 54.0 44.8 44.7 46.9 47.2 47.7 55.4 49.6 54.6 50.9 52.8 50.0 67.5 50.8 55.5 51.0
Ours 1k 85.0 64.0 66.6 90.2 75.2 74.7 78.1 97.2 77.1 77.6 80.1 86.6 77.5 76.5 77.9 77.4 63.1 70.5 77.5
Ours 1k+ 78.7 62.5 68.4 97.5 67.9 84.0 74.3 99.6 78.2 83.7 94.5 97.1 88.9 89.6 81.2 90.9 77.6 83.7 83.2
Ours 10k 85.7 65.5 68.1 90.5 74.7 75.8 78.4 97.7 77.8 78.1 81.2 87.1 77.2 76.4 78.2 76.4 65.0 72.2 78.1
Ours 10k+ 82.8 67.4 70.7 98.4 71.9 85.4 77.3 99.7 80.2 85.8 95.9 98.2 91.1 89.9 83.8 90.1 79.4 85.5 85.2
Table 5: Comparison with SoTA methods (AUC) on images post-processed by random cropping, resizing and compression.
5 Comparison with the state-of-the-art
In this Section we perform an extensive comparison with SoTA methods in different scenarios. We consider four versions of our approach, with 1,000+1,000 (1k) or 10,000+10,000 (10k) real+fake images in the reference set, and with (+) or without compressed/resized images for augmentation. To ensure a fair comparison, we only include SoTA methods with code and/or pre-trained models publicly available online. They are listed in Tab. 3 and described in the supplementary material.
Generalization analysis. In Tab. 4 we show the results in terms of AUC on 18 generative models. For each method, the results obtained on the same dataset used for training are in light gray, since the ID scenario is of little interest for our analysis. The items in bold, instead, highlight the best OOD performance for each dataset, considering a margin of 1%. We can observe that methods trained on one GAN dataset generally perform well on other datasets of the same family but fare much worse on DM datasets. Then, the roles change for methods trained on DM datasets. This is unsurprising, as generators of the same family share architectural details that leave similar traces on the generated images. No SoTA method performs uniformly well on all datasets. In contrast, the proposed lightweight CLIP-based detector consistently delivers strong performance. The version with 10k+10k reference images, without augmentation, outperforms the best competitor by +6.8% in terms of average AUC. For GAN-based generators, the proposed CLIP-based detectors keep providing the best performance, generally much better than SoTA methods, except for the Liu method which works almost at the same level. However, the performance of this method, as well as many other methods, decreases catastrophically when considering synthetic images from commercial tools. This is the most realistic and interesting scenario, with images of unknown origin and no prior information on the possible generation process. In this situation, most methods provide unreliable decisions. Corvi’s method works surprisingly well on some datasets, perhaps generated by diffusion models similar to Latent. However, only the proposed CLIP-based approach provides good stable performance across all cases. On commercial tools, the versions with resized and recompressed images in the reference set prove especially strong.
Robustness to perturbations. The observed trend with respect to unknown models is further accentuated when the images undergo post-processing, as shown in Tab. 5. These impairments attenuate forensic traces, to the point that most SoTA methods become essentially useless, performing no better than random chance, particularly on unknown commercial models. CLIP-based detectors, instead, keep providing a good performance.
6 Beyond low-level forensic traces
As previously mentioned, conventional detectors often experience significant performance degradation when images are resized or compressed, as the subtle forensic traces they depend on are significantly reduced. In contrast, the proposed CLIP-based detector keeps working well under these conditions, suggesting that it relies on higher-level semantic features. To further investigate this hypothesis, we now conduct some targeted experiments.
Removing/inserting low-level traces in images. We take synthetic images generated by Stable Diffusion 1.4, Stable Diffusion 2.0 [63], Stable Diffusion XL [59]) and downsample them by 1/4. Fig. 5 shows, on the left, the high- and low-resolution images (top) along with their spectra (bottom) for Stable XL. After reducing the resolution, the peaks in the Fourier domain completely disappeared, because they were caused by the aliasing effect of the oversampling filters, which can be removed by appropriate decimation. In a complementary experiment, we take real images from the RAISE dataset and pass them through the autoencoders used by Stable Diffusion XL. This process does not change the visual appearance of the images (Fig. 5 top-right) but inserts low-level forensic traces in their spectra (bottom) that resemble those of the generated images.
In Table 6 we quantify the effects of such attacks on 1,000 real and 1,000 synthetic images, reporting the AUC for our method and the best competitor before the attack (first group of columns) after removing traces from synthetic images (second group) and after adding them to real images (third group). Before the attack, all detectors work quite well, with the method proposed by Corvi et al., based on low-level features, reaching perfect discrimination. After the attacks, however, the performance of Corvi et al. degrades dramatically, while the proposed method suffers only a very limited loss. In summary, the CLIP-based detector appears to withstand various types of image degradations easily, both innocent and malicious. The latter case is particularly relevant: attacks can be carried out with the very aim of making everything appear as synthetic, misleading current forensic detectors based on low-level traces, and undermining public trust in forensic analyses.
Fusion. Based on the above results, the proposed CLIP-based detector appears not to rely on the same low-level traces exploited by most of the current detectors. This is further supported by the scatter plot of Fig.6 where each point gives the scores of the CLIP-based and Corvi detectors for a given image. The two sets of scores are almost orthogonal, as if they depended on uncorrelated features, a property that paves the way for appropriate fusion strategies. We implemented a simple decision rule where the image is declared real only if both detectors agree on this choice. Tab.7 reports AUC and Accuracy results for Corvi, the proposed method (10k and 10k+ versions), and their fusion. Results are given as averages over the three families of generators: GAN, Diffusion and Commercial Tools. The fusion ensures a further boost in performance over the proposed CLIP-based detector, both in terms of AUC (+3.6%) and Accuracy (+7.4%). A smaller improvement is observed when images are resized/compressed, arguably because low-level traces are more compromised in this case.